What is the equation of the directrix of a parabola y^2=6x
Answer by DSMLMD(16)

To find the equation of the directrix of parabola, first we must find p value from the parabola equation. Next, we substitute the p value to directrix formula x = -p and change the p value position to left side become x + p = 0.


Parabola equation:


y^2 = 6x
y^2 = 4(3/2)x


Directrix line:


x = -p
x = -(3/2)
x + (3/2) = 0
2x + 3 = 0 (multiplied 2 to both side)


So the equation of the directrix or a parabola y^2 = 6x is 2x + 3 = 0